# Value Alignment of Social Media Ranking Algorithms

### v1

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the CHI Conference on Human Factors in Computing Systems; April 26–May 1, 2025; Yokohama, Japanbooktitle: Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26–May 1, 2025, Yokohama, Japan, USAisbn: 978-1-4503-XXXX-X/25/05ccs: Human-centered computing Collaborative and social computing systems and tools
@@ 1. Introduction
To value align feed algorithms, in this paper, we develop an approach that operationalizes a structured value system—one that is grounded in the empirical behavioral literature and accounts for trade-offs between values—into social media ranking algorithms. Specifically, we adopt Schwartz’s theory of Basic Human Values (Schwartz et al., 2012), which has been widely used and replicated in Cultural Psychology. The modern Schwartz theory articulates 19 conceptually distinct values, each of which is placed into a circumplex model to articulate values that are complementary or in tension with each other. For example, Achievement captures an individualistic focus, whereas Caring represents an opposite and complementary societal focus. The Cultural Psychology literature argues that these values are the basic motivations that drive behavior universally across cultures (Schwartz, 2012b), and can be thought of as the building blocks of other, more specific, values.
@@
We contribute an algorithm for classifying expressions of Schwartz’s Basic Human Values in social media posts, then use these models to rerank social media feeds. We draw on the insight that the value definitions defined by social scientists and humanities scholars are precise enough constructs (Schmitt et al., 1993) for large language models (LLMs) to annotate as accurately as experts do (Jia et al., 2024; Halterman and Keith, 2024; Tai et al., 2024; Dai et al., 2023; Ziems et al., 2024). Leveraging this insight, we develop a technical approach to classify social media posts using Schwartz’s values, then integrate those classifications with users’ value priorities in order to rerank their feed to be aligned with their desired values.
@@
But even if we technically rerank feeds to reflect different values, we need to understand if the resulting value-aligned feeds resonate with users in ways that they can recognize. Recognizablility acts as a proxy for whether our value alignment pipeline, from elicitation to ranking, is translating users’ values into feeds that are perceptibly different and coherent. We demonstrate through two controlled experiments that users can distinguish between value-aligned and engagement-aligned rankings of their own social media feeds, and that this recognizability weakens as users apply more complicated combinations of values. In both experiments, we rank users’ own Twitter/X feeds by the values that they deem important, and present them along the default engagement feed in a blinded selection task.
@@
In the first experiment, we derive users’ self-reported value priorities from a value questionnaire that they fill out. In the second, we offer users value controls that they can directly manipulate and whose effects on their feed they can observe in real time. We take the configured values and apply them to users’ own Twitter feeds, then task users with identifying which is the value-aligned feed and which is the original engagement-aligned feed. Finally, through analyzing 228 re-ranked feeds and conducting think-aloud studies with thirteen participants, we find that users often configure controls that align with their personal values, leading to value-aligned feeds that diverge substantially from engagement-driven feeds (mean Kendall’s $\tau$ of $0.06$).
@@
    Empirical and qualitative insights into value selections: We analyze what mechanisms drive users’ value selections for their feed and how they navigate trade-offs in values.
Ultimately, this work advocates that social media feeds need neither disregard values nor selectively prioritize only a few, but can consider a broad space of value tradeoffs and empower users, platforms, and society to better navigate them.
@@ 2.1. Incorporating human values in ranking algorithms
At the moment, social media platforms primarily utilize engagement signals, such as number of clicks, time spent on an item, or number of comments, to inform feed ranking (Backstrom, 2016; Narayanan, 2023; Bernstein et al., 2023; Cunningham et al., 2024; Milli et al., 2025). Compared to chronological feeds, engagement-based feeds increase user retention (Cunningham et al., 2024; Guess et al., 2023a; Bandy and Lazovich, 2023). However, focusing solely on engagement signals can lead to detrimental social outcomes, including the marginalization of perspectives (Harris et al., 2023; Delmonaco et al., 2024), proliferation of misinformation (Valenzuela et al., 2019; Jahanbakhsh et al., 2021; Hao, 2021), political polarization (Rathje et al., 2021; Kubin and von Sikorski, 2021), and promotion of extremist content (O’Callaghan et al., 2015). To address these issues, recent work has argued that the signals we incorporate into ranking algorithms should be expanded to also take into account broader societal values (Bernstein et al., 2023). While this commitment may lead to tradeoffs with engagement, early evidence did not find measurable engagement tradeoffs when integrating a societal value (Jia et al., 2024).
@@
The challenging questions are: (1) which social values should be embedded in algorithms, and (2) how do we do so? In the social media setting, many prior works have focused on political aspects of this problem, motivated by the pressing concerns about these technologies’ impact on democracy (Jia et al., 2024; Keegan, 2016; Babaei et al., 2018; Nelimarkka et al., 2019; Munson and Resnick, 2010; Munson et al., 2013; Guess et al., 2023b; Nyhan et al., 2023; Milli et al., 2025). There have also been concerted efforts to reduce the amount of misinformation on feeds (Jahanbakhsh et al., 2021; Bode and Vraga, 2018; Karduni et al., 2019; Vraga and Bode, 2017). However, focusing only on misinformation or political content covers only a small fraction of what is shown on social media. Perhaps closest to our work is Alexandria (Kolluri et al., 2025), a library of values spanning six value systems (e.g., Rokeach and Maslow). While Alexandria offers broad coverage, it does not encode relationships of tension or complementarity among values, a limitation that matters when algorithms must inevitably prioritize some values at the expense of others. Other efforts to advocate for a broader set of values have remained either at the theoretical level (Bernstein et al., 2023; Stray et al., 2024; Milli et al., 2021). This leaves questions about how we might integrate a design space of values—one with trade-offs—into algorithms in practice.
@@ 2.2. Providing end-user control on social media
The approach that we detail in this paper can be executed at the central platform level (e.g., Facebook, Twitter, Instagram, TikTok), in a decentralized way by individual communities (e.g., if Mastodon added ranked feeds), or by end users themselves (e.g., Bluesky). As our approach allows end users themselves to specify values, in this section, we review how and why increased community and end user control of feed algorithms can be beneficial.
@@
In an age when much of the information consumption happens on social media by choice or accident, the platforms as the authorities that govern the flow of information wield significant influence over public discourse, shaping narratives and potentially narrowing the diversity of viewpoints. This concentration of power has raised concerns, both in the academic sphere and among the general public, about bias, censorship, or the ad-hoc development and inconsistent enforcement of rules and policies (Saltz et al., 2021; Díaz and Hecht-Felella, 2021; Schneider, 2024). These concerns are grounded in accounts of platforms removing content that arguably posed little to no harm (Sankin, 2017; Fac, 2018; Alimardani, 2022), disproportionately policing speech from minority groups (Sankin, 2017; Jan and Dwoskin, 2017; Sanchez et al., 2024), selectively intervening in certain high-profile cases by bending or outright breaking community standards (Biddle, 2022; Facebook, 2024), and pursuing inaction in others with grave ramifications (Mozur, 2018; Biddle, 2022), among other troubling practices.
@@
Even if we do not view platforms as unfit governors (Klonick, 2017) of speech, we must still seriously weigh the benefits of granting end users more control over their feed experiences. Centralized decisions about what content to show or suppress can fail to account for the diverse needs and desires of each user or community. For example, while platforms typically down-rank misinformation, prior work has shown that some users prefer to avoid unreliable content entirely, while others want to see it to understand what misinformation their friends and family see in order to talk to them about it (Jahanbakhsh et al., 2022a; Malhotra, 2020; Rader and Gray, 2015). As another example, minority groups, who use social media to call out racism and recount the language used against them, report being found in violation of centralized guidelines, as their posts are misconstrued as hate speech (Jan and Dwoskin, 2017; Lee et al., 2024).
@@
The centralized control of information on platforms is primarily executed through content curation and moderation, much of which is handled algorithmically. In the absence of end-user controls explicitly built into the process, users resort to finding ways around the algorithm, albeit with uncertain efficacy, by developing “folk theories” that rationalize how their inputs translate into outputs (Eslami et al., 2015). For instance, to avoid seeing content they do not like, they might “like other things such that these show up less” (Jahanbakhsh et al., 2022a; Eslami et al., 2016).
@@
With the growing awareness about the shortcomings of one-size-fits-all approaches and the recognition that users do desire control, there have been endeavors to give users more agency over their information space. Notable examples in the industry include Mastodon, a decentralized social network where users can join servers with varying curation and moderation policies, and Bluesky, which allows users to subscribe to different feed algorithms or moderation modules (Team, 2024). Tools such as Skyfeed [^1] https://skyfeed.app/ and Graze [^2] https://www.graze.social/ utilize Bluesky’s API, offering users the ability to author custom feeds. A body of academic work has also concerned itself with pluralist models of governance that range from delegating the power of decision-making to community juries (Fan and Zhang, 2020; Micek and Solovey, 2024), to empowering the individual user (Jhaver and Zhang, 2023; Burrell et al., 2019). These have been instantiated in the form of filters or extensions on top of existing platforms (Bhargava et al., 2019; Jahanbakhsh et al., 2022b; Jahanbakhsh and Karger, 2024; Jhaver et al., 2023) as well as separate social media platforms (Jahanbakhsh et al., 2022a). Another line of work has proposed new methods for personalizing the feed experience using AI tools (Feng et al., 2024a; Jahanbakhsh et al., 2023; Wang et al., 2025).
@@
Building on this literature, we design an interface that allow users to specify their value preferences and directly reconfigure their feeds. While prior work focused on empowering interventions in more targeted use cases, such as reducing misinformation (Jahanbakhsh et al., 2022a) and combating harassment (Jhaver et al., 2023), or only provided a limited set of filters (Bhargava et al., 2019), we present a more comprehensive set of controls. By focusing on the basic values present in the feed, and incorporating those into the algorithm, our controls can be applied to the broad range of content that is available on social media.
@@ 3. Value Aligning Social Media Feeds with Schwartz’s Basic Values
To enable our social media feeds to become responsive to values and their tradeoffs, we require a complete value system: a set of values that anchor the basic tensions at stake. Of course, there are many possible values: a recent paper articulated over two hundred thousand values, rights, and duties (Sorensen et al., 2024). Instead of seeking an articulation of every single value, we instead advocate for structural coverage in the sense that the set of values encompasses the main axes of differentiation in the design space, and that most specific values can be thought of as specializations of those in our set.
In our work, we use Schwartz’s theory of basic values, which provides a set of universally recognized concepts, for operationalizing these constructs in social media content. Schwartz’s system has been applied not only for looking at individuals’ values but also for analyzing values in text, including news articles and arguments (Bardi et al., 2008; Kiesel et al., 2022; Borenstein et al., 2024). While Schwartz’s theory is not the only lens through which values can be investigated, it does provide a comprehensive value system that has been extensively validated across different populations (Schwartz and Bilsky, 1990; Schwartz and Butenko, 2014; Schwartz et al., 2017). Of course, other complete value sets may also satisfy these criteria (e.g., Moral Foundations Theory (Graham et al., 2013)). Our method aims to generalize to any value system that satisfies these criteria.
@@ 3.1.1. Classifying values
Next, to allow users to re-rank their feeds, we need an automated way to code value expression scores (Piccardi et al., 2024). As prior work has demonstrated (Jia et al., 2024), LLMs can apply constructs from social sciences for coding social media content. In our work, we use few-shot prompting with GPT-4o to code for Schwartz’s values in tweets. GPT-4o and similar multi-modal models (e.g., GPT-4o-mini, Gemini, Llama) are an attractive option for this task because their training has exposed them to a broad range of cultural, political, and topic themes, and their ability to take in images as well as text fits well with the social media context. As input, we give the model the text of the tweet, accompanying images, and embedded links along with their descriptions. In addition to two examples that were manually coded by the researchers, we provide definitions of the 19 values taken from Schwartz’s theory, and the tweet itself, as input to the model. Since Schwartz argues that values are in fact not independent from each other and share the same (or conflicting) underlying motivations, we included all 19 values in one prompt rather than building individual classifiers. As output, the model returns ratings for all 19 values in JSON format. See the Appendix for prompt details and examples of coded tweets.

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791281isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Collaborative and social computing systems and tools
@@
Diagram illustrating the proposed value-aligned feed ranking method. Diagram illustrating the proposed value-aligned feed ranking method. Social media posts are classified by a large language model into Schwartz’s values, combined with user-specified weights, and ranked accordingly. The figure shows an example where weights for “Caring,” “Tolerance,” and “Tradition” are specified, demonstrating how these weights alter the ranking.
@@ 1. Introduction
To value align feed algorithms, in this paper, we develop an approach that operationalizes a structured value system — one that is grounded in the empirical behavioral literature and accounts for trade-offs between values — into social media ranking algorithms. Specifically, we adopt Schwartz’s theory of Basic Human Values (Schwartz et al., 2012), which has been widely used and replicated in Cultural Psychology. The modern Schwartz theory articulates 19 conceptually distinct values, each of which is placed into a circumplex model to articulate values that are complementary or in tension with each other. For example, Achievement captures an individualistic focus, whereas Caring represents an opposite and complementary societal focus. The Cultural Psychology literature argues that these values are the basic motivations that drive behavior universally across cultures (Schwartz, 2012), and can be thought of as the building blocks of other, more specific, values.
@@
We contribute an algorithm for classifying expressions of Schwartz’s Basic Human Values in social media posts, then use these models to rerank social media feeds (Fig. 1). We draw on the insight that the value definitions defined by social scientists and humanities scholars are precise enough constructs (Schmitt et al., 1993) for large language models (LLMs) to annotate as accurately as experts do (Jia et al., 2024; Halterman and Keith, 2025; Tai et al., 2024; Dai et al., 2023; Ziems et al., 2024). Leveraging this insight, we develop a technical approach to classify social media posts using Schwartz’s values, then integrate those classifications with users’ value priorities in order to rerank their feed to be aligned with their desired values.
But even if we technically rerank feeds to reflect different values, we need some assurance that users’ value specifications are preserved in a meaningful way rather than being collapsed into noise. When multiple values interact, and often stand in tension, recognizability of the values in the resulting feed becomes a necessary condition for establishing that the feed meaningfully reflects those specifications. If users can perceive a coherent relationship between the values they articulate and the feed they receive, then the pipeline from elicitation to ranking is at least preserving a meaningful input–output mapping.
@@
We demonstrate through two controlled experiments that users can distinguish between value-aligned and engagement-aligned rankings of their own social media feeds, and that this recognizability weakens as users apply more complicated combinations of values. In both experiments, we rank users’ own Twitter/X feeds by the values that they deem important, and present them along the default engagement feed in a blinded selection task.
@@
In the first experiment, we derive users’ self-reported value priorities from a value questionnaire that they fill out. In the second, we offer users value controls that they can directly manipulate and whose effects on their feed they can observe in real time. We take the configured values and apply them to users’ own Twitter feeds, then task users with identifying which is the value-aligned feed and which is the original engagement-aligned feed. Finally, through analyzing 188 re-ranked feeds, we find that users often configure controls that align with their personal values, leading to value-aligned feeds that diverge substantially from engagement-driven feeds (mean Kendall’s $\tau$ of $0.06$).
@@
    Empirical insights into value selections: We analyze what mechanisms drive users’ value selections and how these selections impact feed rankings.
Ultimately, this work advocates that social media feeds need neither disregard values nor selectively prioritize only a few, but can consider a broad space of value tradeoffs and empower users, platforms, and society to better navigate them. [^1] Code and project website available at https://stanfordhci.github.io/FeedValueAlignment/.
@@ 2.1. Incorporating human values in ranking algorithms
At the moment, social media platforms primarily utilize engagement signals, such as number of clicks, time spent on an item, or number of comments, to inform feed ranking (Backstrom, 2016; Narayanan, 2023; Bernstein et al., 2023; Cunningham et al., 2024; Milli et al., 2025). Compared to chronological feeds, engagement-based feeds increase user retention (Cunningham et al., 2024; Guess et al., 2023a; Bandy and Lazovich, 2023). However, focusing solely on engagement signals can lead to detrimental social outcomes, including the marginalization of perspectives (Harris et al., 2023; Delmonaco et al., 2024), proliferation of misinformation (Valenzuela et al., 2019; Jahanbakhsh et al., 2021; Hao, 2021), political polarization (Rathje et al., 2021; Kubin and von Sikorski, 2021), and promotion of extremist content (O’Callaghan et al., 2015). It is not only what content users see, but also the ranking of the information can impact users’ opinion formation and even influence downstream actions, such as swaying voting preferences (Epstein and Robertson, 2015; Draws et al., 2021a; Chan et al., 2025). To address these issues, recent work has argued that the signals we incorporate into ranking algorithms should be expanded to also take into account broader societal values (Bernstein et al., 2023). While this commitment may lead to tradeoffs with engagement, early evidence did not find measurable engagement tradeoffs when downranking content that exhibited anti-democratic attitudes (Jia et al., 2024).
@@
The challenging questions are: (1) which social values should be embedded in algorithms, and (2) how do we do so? In the social media setting, many prior works have focused on political aspects of this problem, motivated by the pressing concerns about these technologies’ impact on democracy (Jia et al., 2024; Keegan, 2016; Babaei et al., 2018; Nelimarkka et al., 2019; Munson and Resnick, 2010; Munson et al., 2013; Guess et al., 2023b; Nyhan et al., 2023; Milli et al., 2025). There have also been concerted efforts to reduce the amount of misinformation on feeds (Jahanbakhsh et al., 2021; Bode and Vraga, 2018; Karduni et al., 2019; Vraga and Bode, 2017). However, focusing only on misinformation or political content covers only a small fraction of what is shown on social media. Perhaps closest to our work is Alexandria (Kolluri et al., 2025), a library of values spanning six value systems (e.g., Rokeach and Maslow). While Alexandria offers broad coverage, it does not encode relationships of tension or complementarity among values; this limitation is important as algorithms must inevitably prioritize some values at the expense of others. In contrast, Schwartz’s value system provides a theoretical structure for reasoning over trade-offs between values. Furthermore, our work validates that our method using direct user elicitation and LLM value labels produces recognizably value-aligned feed and additionally provides empirical insights into how this value alignment reshapes users’ feeds. Finally, our studies reveal that maintaining recognizability becomes difficult when users adjust multiple values simultaneously, though this decline plateaus rather than continuing to degrade. This finding has concrete implications for designing value elicitation mechanisms. Outside of Alexandria, other efforts to advocate for a broader set of values have remained at the theoretical level (Bernstein et al., 2023; Stray et al., 2024; Milli et al., 2021; Rieger et al., 2024). This leaves questions about how we might integrate a design space of values — one with trade-offs — into algorithms in practice.
@@ 2.2. Providing end-user control on social media
The approach that we detail in this paper can be executed at the central platform level (e.g., Facebook, Twitter, Instagram, and TikTok), in a decentralized way by individual communities (e.g., if Mastodon added ranked feeds), or by end users themselves (e.g., Bluesky). As our approach allows end users themselves to specify values, in this section, we review how and why increased community and end user control of feed algorithms can be beneficial.
@@
In an age when much of the information consumption happens on social media by choice or accident, the platforms as the authorities that govern the flow of information wield significant influence over public discourse, shaping narratives and potentially narrowing the diversity of viewpoints (Draws et al., 2021b; Bozdag and Van Den Hoven, 2015; Kitchens et al., 2020). This concentration of power has raised concerns, both in the academic sphere and among the general public, about bias, censorship, or the ad-hoc development and inconsistent enforcement of rules and policies (Saltz et al., 2021; Díaz and Hecht-Felella, 2021; Schneider, 2024). These concerns are grounded in accounts of platforms removing content that arguably posed little to no harm (A. Sankin (2017); 34; M. Alimardani (2022)), disproportionately policing speech from minority groups (Sankin, 2017; Jan and Dwoskin, 2017; Sanchez et al., 2024), selectively intervening in certain high-profile cases by bending or outright breaking community standards (Biddle, 2022; Facebook, 2024), and pursuing inaction in others with grave ramifications (Mozur, 2018; Biddle, 2022), among other troubling practices.
@@
Even if we do not view platforms as unfit governors (Klonick, 2017) of speech, we must still seriously weigh the benefits of granting end users more control over their feed experiences. Centralized decisions about what content to show or suppress can fail to account for the diverse needs and desires of each user or community. For example, while platforms typically down-rank misinformation, prior work has shown that some users prefer to avoid unreliable content entirely, while others want to see it to understand what misinformation their friends and family see in order to talk to them about it (Jahanbakhsh et al., 2022b; Malhotra, 2020; Rader and Gray, 2015). As another example, minority groups, who use social media to call out racism and recount the language used against them, report being found in violation of centralized guidelines, as their posts are misconstrued as hate speech (Jan and Dwoskin, 2017; Lee et al., 2024).
@@
The centralized control of information on platforms is primarily executed through content curation and moderation, much of which is handled algorithmically. In the absence of end-user controls explicitly built into the process, users resort to finding ways around the algorithm, albeit with uncertain efficacy, by developing “folk theories” that rationalize how their inputs translate into outputs (Eslami et al., 2015). For instance, to avoid seeing content they do not like, they might “like other things such that these show up less” (Jahanbakhsh et al., 2022b; Eslami et al., 2016).
@@
With the growing awareness about the shortcomings of one-size-fits-all approaches and the recognition that users do desire control, there have been endeavors to give users more agency over their information space. Notable examples in the industry include Mastodon, a decentralized social network where users can join servers with varying curation and moderation policies, and Bluesky, which allows users to subscribe to different feed algorithms or moderation modules (Team, 2024). Tools such as Skyfeed [^2] https://skyfeed.app/ and Graze [^3] https://www.graze.social/ utilize Bluesky’s API, offering users the ability to author custom feeds. A body of academic work has also concerned itself with pluralist models of governance that range from delegating the power of decision-making to community juries (Fan and Zhang, 2020; Micek and Solovey, 2024), to empowering the individual user (Jhaver and Zhang, 2023; Burrell et al., 2019). These have been instantiated in the form of filters or extensions on top of existing platforms (Bhargava et al., 2019; Jahanbakhsh et al., 2022a; Jahanbakhsh and Karger, 2024; Jhaver et al., 2023) as well as separate social media platforms (Jahanbakhsh et al., 2022b). Another line of work has proposed new methods for personalizing the feed experience using AI tools (Feng et al., 2024a; Jahanbakhsh et al., 2023; Wang et al., 2025).
@@
Building on this literature, we design an interface that allow users to specify their value preferences and directly reconfigure their feeds. While prior work focused on empowering interventions in more targeted use cases, such as reducing misinformation (Jahanbakhsh et al., 2022b) and combating harassment (Jhaver et al., 2023), or only provided a limited set of filters (Bhargava et al., 2019), we present a more comprehensive set of controls. By focusing on the basic values present in the feed, and incorporating those into the algorithm, our controls can be applied to the broad range of content that is available on social media.
@@ 3. Value Aligning Social Media Feeds with Schwartz’s Basic Values
To enable our social media feeds to become responsive to values and their tradeoffs, we require a complete value system: a set of values that anchor the basic tensions at stake. Of course, there are many possible values (i.e., Sorensen et al. (2024) articulated over two hundred thousand values, rights, and duties). Instead of seeking an articulation of every single value, we instead advocate for structural coverage in the sense that the set of values encompasses the main axes of differentiation in the design space, and that most specific values can be thought of as specializations of those in our set.
In our work, we use Schwartz’s theory of basic values, which provides a set of universally recognized concepts, for operationalizing these constructs in social media content. Schwartz’s system has been applied not only for looking at individuals’ values but also for analyzing values in text, including news articles and arguments (Bardi et al., 2008; Kiesel et al., 2022; Borenstein et al., 2025).
According to Schwartz, values transcend specific situations and serve as standards or criteria for what is considered good or bad. Trade-offs among these broad motivational principles guide attitudes and behaviors across a wide range of contexts. For example, values such as obedience or honesty may be relevant whether one is interacting with friends, colleagues, or strangers (Schwartz, 2012, 1992, 2013). Schwartz further shows that individuals’ general value priorities, elicited in a domain-independent way, predict patterns of behavior across diverse contexts, from interpersonal cooperation to voting behavior to intergroup contact (Schwartz, 2013). This trans-situational quality distinguishes values from norms or attitudes, which tend to be context-specific (Schwartz, 2012). Because Schwartz’s values operate at this general level and have been extensively validated across diverse contexts, they can be elicited in a domain-general form and then operationalized in specific contexts such as social media consumption.
While Schwartz’s theory is not the only lens through which values can be investigated, it does provide a comprehensive value system that has been extensively validated across different populations (Schwartz and Bilsky, 1990; Schwartz and Butenko, 2014; Schwartz et al., 2017). Of course, other complete value sets may also satisfy these criteria (e.g., Moral Foundations Theory (Graham et al., 2013)). Our method aims to generalize to any value system that satisfies these criteria.
@@ 3.1. Measuring value expressions in social media posts
Circular diagram of Schwartz’s 19 Basic Human Values. Circular diagram of Schwartz’s 19 Basic Human Values. Values are organized on a circumplex, where adjacent values share similar motivations, and opposing values are in tension. The four broader groups—self-transcendence, conservation, self-enhancement, and openness to change—are highlighted
@@ 3.1.1. Classifying values
Next, to allow users to re-rank their feeds, we need an automated way to code value expression scores (Piccardi et al., 2024). As prior work has demonstrated (Jia et al., 2024), LLMs can apply constructs from social sciences for coding social media content. In our work, we use few-shot prompting with GPT-4o to code for Schwartz’s values in tweets. GPT-4o and similar multi-modal models (e.g., GPT-4o-mini, Gemini, and Llama) are an attractive option for this task because their training has exposed them to a broad range of cultural, political, and topic themes, and their ability to take in images as well as text fits well with the social media context. As input, we give the model the text of the tweet, accompanying images, and embedded links along with their descriptions. In addition to two examples that were manually coded by the researchers, we provide definitions of the 19 values taken from Schwartz’s theory, and the tweet itself, as input to the model. Since Schwartz argues that values are in fact not independent from each other and share the same (or conflicting) underlying motivations, we included all 19 values in one prompt rather than building individual classifiers. As output, the model returns ratings for all 19 values in JSON format. See the Appendix for prompt details and examples of coded tweets.